Solve equation for x.
1-(2/x)-(15/x^2)=0

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Solve equation for x.
1-(2/x)-(15/x^2)=0
---------------
Multiply thru by x^2 to get:
x^2 - 2x - 15 = 0
Factor
(x-5)(x+3) = 0
x = 5 or x = -3
==========================
Cheers,
Stan H.

Answer by checkley77(12844)

1-(2/x)-(15/x^2)=0 [multiply 1 by x^2 & the 2/x term by x & add them to the 15/x^2 term as a numerator for the x^2 term.
(x^2-2x-15)/x^2=0
x^2-2x-15=0
(x-5)(x+3)=0
x-5=0 ans.
x+3=0
x=-3 ans.
Proof:
1-(2/5)-(15/5^2)=0
1-.4-15/25=0
1-.4-.6=0
0=0
1-(2/-3)-(15/3^2)=0
1+2/3-15/9=0
1+.667-1.667=0
0=0